OpenStudy (anonymous):

sodium reacts with chlorine gas to form sodium chloride. 2Na+CL2--> 2NaCl what mass of chlorine gas will react with 92.0 g of sodium

OpenStudy (anonymous):

Not entirely sure if its correct by I get 141.80 grams Cl2. \[92g Na \times \frac{ 1mol }{ 23g }=4molNa \times \frac{ 1mol Cl _{2} }{ 2molNa}=2mol Cl _{2}\times \frac{ 70.9g }{ 1mole}=141.80g Cl _{2}\]